OpenStudy (idkanything):

A certain computer can perform 10^5 calculations per second. How many calculations can it perform in 10 seconds? I used scientific notation but i think it is wrong @tshimp0629

OpenStudy (idkanything):

@3mar

OpenStudy (idkanything):

@KJSaif

OpenStudy (jango_in_dtown):

1sec= 10^5 calculations 10sec=10*10^5 calculations= ???

OpenStudy (idkanything):

1,000,000??

OpenStudy (jango_in_dtown):

correct

OpenStudy (idkanything):

so thats the answer?

KjSaif (kjsaif):

Your right

OpenStudy (idkanything):

oh ok thanks

OpenStudy (3mar):

So sorry, I was not here!